Check whether the series sum_(n=1)^(oo)(nx)/(n^(4)+x^(3)) x in 0 alpha is uniformly convergent or not

1
Expert's answer
2022-03-28T18:15:40-0400

ANSWER The series "\sum_{n=1}^{\infty}\frac{nx}{n^{4}+x^{3}}" is uniformly convergent on "\left [ 0, \alpha \right ]" .

EXPLANATION.

Since a) "0\leq nx\leq n \alpha"

b) "n^{4}+x^{3}\geq n^{4}"

for "x\in[0,\alpha]" and "n\geq1" , then

"0\leq\frac{nx}{n^{4}+^x{3}}\leq\frac{n\alpha}{n^{4}}=\frac{\alpha}{n^{3}}" .

The series "\sum_{n=1}^{\infty}\frac{\alpha}{n^{3}}=\alpha\sum_{n=1}^{\infty }\frac{1}{n^{3}}" is convergent , because the series "\sum_{n=1}^{\infty }\frac{1}{n^{3}}" is a "p"-series for "p=3" .

Therefore, by the Weierstrass M-Test the series "\sum_{n=1}^{\infty}\frac{nx}{n^{4}+x^{3}}" is uniformly convergent on "\left [ 0, \alpha \right ]" .